A heartbroken fellow throws his engagement ring from the roof of a 20.0-m-tall building at an angle of  above the horizontal and with an initial speed of . Ignore air resistance and calculate:


a.     the maximum height above the roof of the building that the ring reaches;

b.    the speed of the ring before it reaches the ground;

c.     the horizontal range from the base of the building to the point where the ring strikes the ground.



1
Expert's answer
2021-09-29T14:08:03-0400

Let angle "\alpha=30\degree" , initial speed "v_0=30" m/s


a.

"a_y=\frac{dv_y}{dt}=-g"

"v_y=-gt+v_0sin\alpha"

"h=-gt^2/2+v_0tsin\alpha"

"h'(t)=-gt+v_0sin\alpha=0"

"t=v_0sin\alpha /g=30\cdot 0.5/9.8=1.53" s

"h_{max}=30\cdot 1.53 \cdot 0.5-9.8\cdot 1.53^2/2=11.5" m


b.

"h=-gt^2/2+v_0tsin\alpha+h_0=0"

"h_0=20" m

"-4.9t^2+15t+20=0"

"t=\frac{15+\sqrt{225+80\cdot 4.9}}{9.8}=4.07" s

"v_y=-9.8\cdot 4.07+15=-24.89" m/s

"v_x=v_0cos\alpha=30\sqrt{3}/2=25.98" m/s

"v=\sqrt{v_x^2+v_y^2}=\sqrt{24.89^2+25.98^2}=35.98" m/s


c.

"s=v_xt=v_0tcos\alpha=25.98\cdot 4.07=105.74" m